show how to solve 3x^2 + 10x = 8 by factoring
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

show how to solve 3x^2 + 10x = 8 by factoring
3x^2 + 10x - 8 = 0
Step 1, factor it.
Factoring is trial and error, but you get better with practice.
Having someone do it for you doesn't accomplish anything.

Answer by stanbon(75887)

solve 3x^2 + 10x = 8 by factoring
-----
3x^2 + 10x - 8 = 0
-----
Using the AC Method::
A*C = 3*-8 = -24
B = 10
-----
-2*12 = -24
-2+12 = 10
-----
Rewrite the equation::
3x^2 +12x -2x -8 = 0
-----
3x(x+4)-2(x+4) = 0
-----
(x+4)(3x-2) = 0
-----
x = -4 or x = 2/3
------------------------
Cheers,
Stan H.
--------------